A block is pressed against a vertical wall by a fo
Answer #1

a) When it slides up, the sum of forces in vertical direction would be:

F cos\theta - \muk F sin\theta - W = 0

F = W/ (cos \theta- \muk sin\theta )

F = 55/( cos46 - 0.33 x sin46) = 120.35 N

Hence, F = 120.35 N

b) in this case,

W - F cos\theta - \muk F sin\theta = 0

F = W/ (cos\theta + \muk sin\theta)

F = 55/ (cos46 + 0.33 x sin46) = 59.01 N

Hence, F = 59.01 N